OpenStudy (anonymous):

PLEASE HELPP MEDALLS. What are the roots of -12x^2+11x-3=0

OpenStudy (zzr0ck3r):

you can complete the square or use the quadratic formula

OpenStudy (zzr0ck3r):

do you know how to do either?

OpenStudy (anonymous):

Yeah let me try the quadratic formula and see if i get the right answer

OpenStudy (anonymous):

Yeah...idk how to do it.

OpenStudy (zzr0ck3r):

\(ax^2+bx+c=0 \iff x=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a} \) for you \[a=-12 \\ b=11\\c=-3\]

OpenStudy (anonymous):

Yeah i got that ..but the answer choices have the letter i in them and idk where the i comes from

OpenStudy (anonymous):

oh that is complex numbers used when you get negative square root

OpenStudy (anonymous):

Would the answer be -11 plus or negative i×square root of 15 over 12?

OpenStudy (anonymous):

I meant square root of 23 over 24

OpenStudy (anonymous):

basically \[i = \sqrt{-1}\]

OpenStudy (anonymous):

\[e.g.\sqrt{-11}=\sqrt{11}i\]

OpenStudy (anonymous):

I understand thank you!

OpenStudy (anonymous):

no worries :)